# A novel efficient structure-preserving exponential integrator for Hamiltonian systems

### v1

@@ 2 The new linearly implicit structure-preserving symmetric method
$\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{T}\hat{\nabla}U(x_{n+1},x_{n})=0,$
where $A=QM$, and ^$\hat{\nabla}U(x_{n},x_{n+1})$ is a discretization of the true gradient $\nabla U(x_{n})$.
Proof.
Assuming that the matrix $M$ is not singular and denoting by $V=hA=hQM$ and ^$\hat{\nabla}U=M\breve{\nabla}U$, we get the following equation after replacing $x_{n+1}$ by $e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1})$:
$\begin{split}&\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}{x_{n}}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1})\\ &=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+h{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U\\ &+\frac{1}{2}h^{2}\hat{\nabla}U^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U\\ &+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}-I)\hat{\nabla}U+{\hat{\nabla}U}^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}\hat{\nabla}U\\ &=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U\\ &+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}(e^{V}-I)^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U\\ &+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U+{\breve{\nabla}U}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U\\ &=\frac{1}{2}(x_{n}+\breve{\nabla}U)^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)(x_{n}+\breve{\nabla}U)+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-Me^{V})\breve{\nabla}U\\ &=0.\end{split}$
The last equation follows from the fact that ${e^{V}}^{\mathsf{T}}Me^{V}-M$ and ${e^{V}}^{\mathsf{T}}M-Me^{V}$ are skew symmetric.
If $M$ is singular, we follow the idea in [15], introducing a series of symmetric and nonsingular matrices $M_{\delta}$ which converges to $M$ when $\delta$ approaches zero to complete the proof. ∎
The discretized gradient ^$\hat{\nabla}U(x_{n},x_{n+1})$ can include the discrete gradient obtained by AVF, the discretization given by Kahan’s method as well as a more general discretization by Runge-Kutta method. We denote the discretized gradient obtained via Kahan’s method by $\nabla U_{\text{K}}$, which can be written as
$\nabla U_{\text{K}}=-\frac{1}{2}\nabla U(x_{n})+2\nabla U(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla U(x_{n+1}).$
@@
For a Hamiltonian system (2) with a homogeneous cubic function $U(x)$, EKahan scheme (4) nearly preserves the exact energy and has the following error estimation
$H_{n+1}-H_{n}$ $=U(x_{n+1}-x_{n}),$
@@
Based on Lemma 3, we have
$\begin{split}H_{n+1}-H_{n}&=\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+U(x_{n+1})-U(x_{n})\\ &=U(x_{n+1})-U(x_{n})-(x_{n+1}-x_{n})^{\mathsf{T}}\nabla_{\text{K}}U(x_{n},x_{n+1}).\end{split}$ (15)
For a homogeneous cubic function $U(x)$, we have $U(x)=\frac{1}{6}x^{T}U^{\prime\prime}(x)x$ and $\nabla U_{\text{K}}=\frac{U^{\prime\prime}(x_{n+1})x_{n}}{2}$.
Then, the above energy error (15) can be reformulated as
@@
$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}$
$-\frac{1}{2}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{2}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$+\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}$
$-\frac{1}{3}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{3}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
@@
For Hamiltonian systems (2) with high-order polynomial function $U(x)$, EKahan method can be generalized using the polarization technique presented in [30]. Suppose $U(x)$ is a homogeneous polynomial of order $k+2$, EKahan scheme is given by
$x_{n+k}=e^{khQM}x_{n}+kh\phi(khQM)Q\nabla U_{\text{K}}(x_{n},x_{n+1},\cdots,x_{n+k}),$ (16)
where $\nabla U_{\text{K}}(\cdot,\cdots,\cdot)$ is a symmetric $k+1$-linear polarized function of $\nabla U(x)$, satisfying
$\nabla U_{\text{K}}(x_{n},x_{n+1},\cdots,x_{n+k})=\sum_{\begin{subarray}{c}1\leq m\leq k+1\\ n\leq i_{1}<\cdots<i_{m}\leq n+k\end{subarray}}\frac{(-1)^{k+1-m}}{(k+1)!}\nabla U(x_{i_{1}}+\cdots+x_{i_{m}}).$
In this case, we give the following corollaries to describe EKahan’s properties.
@@
For semilinear system (2) with high-order polynomial function $U(x)$, the EKahan scheme (16) is symmetric and second order if the initial starting points $x_{n+1},\cdots,x_{n+k}$ are given using a at least second-order method.
This corollary is a generalization of Lemma 1 and 2 for Hamiltonian systems (2) with cubic polynomial function $U(x)$ to systems with high-order polynomial functions, and the proof follows similarly.

### v2

@@ 2.1 The one-step EKahan scheme for cubic Hamiltonians
$\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{T}\hat{\nabla}U(x_{n},x_{n+1})=0,$
where $\hat{\nabla}U(x_{n},x_{n+1})$ is a discretization of the true gradient $\nabla U(x_{n})$.
Proofs of this lemma and the subsequent lemmas and corollaries are given in the Appendix.
The discretized gradient $\hat{\nabla}U(x_{n},x_{n+1})$ include the discrete gradient obtained by AVF, the discretization given by Kahan’s method as well as a more general discretization by Runge-Kutta method. We denote the discretized gradient obtained via Kahan’s method by $\nabla_{\mathrm{K}}U$, which can be written as
$\nabla_{\mathrm{K}}U=-\frac{1}{2}\nabla U(x_{n})+2\nabla U(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla U(x_{n+1}).$
For the AVF choice, $\hat{\nabla}U$ is a discrete gradient and Lemma 2 implies exact energy conservation. For the Kahan choice, Lemma 2 will yield the following energy-increment identity.
@@
For a Hamiltonian system (2) with a homogeneous cubic function $U(x)$, the EKahan scheme (3) satisfies the stepwise identity
$H_{n+1}-H_{n}=U(x_{n+1}-x_{n}),$ (16)
@@
Applying Lemma 2 with $\hat{\nabla}U=\nabla_{\mathrm{K}}U$ gives
$H_{n+1}-H_{n}=U(x_{n+1})-U(x_{n})-(x_{n+1}-x_{n})^{\mathsf{T}}\nabla_{\text{K}}U(x_{n},x_{n+1}).$ (17)
For a homogeneous cubic function $U(x)$, we have $U(x)=\frac{1}{6}x^{T}U^{\prime\prime}(x)x$ and $\nabla_{\mathrm{K}}U=\frac{1}{2}U^{\prime\prime}(x_{n+1})x_{n}$. Then (17) can be reformulated as
@@
$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}-\frac{1}{2}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$+\frac{1}{2}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}-\frac{1}{3}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{3}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
@@
Let the potential $U$ be a general cubic polynomial with homogeneous decomposition as
$U(x)=U_{3}(x)+U_{2}(x)+U_{1}(x)+U_{0},$
where $U_{j}$ is homogeneous of degree $j$. Then the EKahan scheme (3) satisfies
$H_{n+1}-H_{n}=U_{3}(x_{n+1}-x_{n}),$
showing that the stepwise energy increment depends only on the cubic component.
Remark 1.
Theorem 2 and Corollary 1 provide a computable description of the one-step energy defect in terms of the numerical increment. This complements standard finite-time estimates obtained indirectly from second-order convergence, and explains the favourable long-time behaviour of EKahan despite not implying exact energy conservation or symplecticity for general nonlinear Hamiltonian systems.
2.2 The multistep EKahan scheme for higher-degree polynomial Hamiltonians
We now extend the one-step EKahan scheme (3) from cubic potentials to polynomial potentials of degree $k+2$ with $k\in\mathbb{Z}^{+}$. The key point is to retain the two features of EKahan: (i) exponential propagation of the linear flow, and (ii) a Kahan-type reversible discretization of the polynomial Hamiltonian part that requires solving only one linear system per time step. To achieve (ii), we portion the nonlinearity by employing the symmetric $(k+1)$-linear polarization introduced in [6], leading to a $k$-step method.
We denote the symmetric $(k+1)$-linear polarization of $\nabla U$ by
$\nabla_{\mathrm{K}}U(x_{n},x_{n+1},\cdots,x_{n+k})=\sum_{\begin{subarray}{c}1\leq m\leq k+1\\ n\leq i_{1}<\cdots<i_{m}\leq n+k\end{subarray}}\frac{(-1)^{k+1-m}}{(k+1)!}\,\nabla U(x_{i_{1}}+\cdots+x_{i_{m}}).$ (18)
Then, we propose the following $k$-step EKahan scheme
$x_{n+k}=e^{khA}x_{n}+kh\phi(khA)Q\nabla_{\mathrm{K}}U(x_{n},x_{n+1},\cdots,x_{n+k}).$ (19)
Lemma 3.
Let $U$ be a polynomial of degree $k+2$. Then for each $n$ there exist a matrix $B_{n}\in\mathbb{R}^{d\times d}$ and a vector $\ell_{n}\in\mathbb{R}^{d}$, depending only on $(x_{n},\dots,x_{n+k-1})$, such that
$\nabla_{\mathrm{K}}U(x_{n},\dots,x_{n+k})=B_{n}x_{n+k}+\ell_{n}.$ (20)
Consequently, (19) is equivalent to the linear system
$\Bigl(I-kh\,\phi(khQM)\,QB_{n}\Bigr)x_{n+k}=e^{khQM}x_{n}+kh\,\phi(khQM)\,Q\ell_{n}.$
In particular, if $U$ is homogeneous, then $\ell_{n}=0$.
We list several physically relevant polynomial potentials and give the corresponding polarized gradient $\nabla_{\mathrm{K}}U$.
  • For the standard lattice Duffing oscillator with on-site potential
    $U(q)=\frac{\beta}{4}\sum_{j=1}^{d}(q^{j})^{4},\qquad\beta>0,$
    the polarization formula (18) (with $k=2$) yields
    $\nabla_{\mathrm{K}}U(q_{n},q_{n+1},q_{n+2})=\beta\,(q_{n}\odot q_{n+1}\odot q_{n+2}),$
    where $\odot$ denotes componentwise product, and the corresponding coefficient matrix and vector in (20) are $B_{n}=\beta\,\mathrm{diag}(q_{n}\odot q_{n+1})$ and $\ell_{n}=0$.
  • For the Fermi-Pasta-Ulam with quartic potential
    $U(q)=\frac{\beta}{4}\sum_{i=1}^{N-1}\bigl(r^{i}(q)\bigr)^{4},\qquad r^{i}(q)=\frac{q^{i+1}-q^{i}}{\Delta x},$
    introduce the strain vector $r(q)=(r^{1}(q),\ldots,r^{N-1}(q))^{\mathsf{T}}$ and $r_{n}=r(q_{n})$. Then polarization formula (18) (with $k=2$) gives
    $\nabla_{\mathrm{K}}U(q_{n},q_{n+1},q_{n+2})=D^{\mathsf{T}}\bigl(\beta\,(r_{n}\odot r_{n+1}\odot r_{n+2})\bigr),$
    where $D$ is the forward difference operator, and the corresponding coefficient matrix and vector in (20) are $B_{n}=D^{\mathsf{T}}\,\mathrm{diag}(w_{n})\,D$ and $\ell_{n}=0$, with $w_{n}=\beta\,(r_{n}\odot r_{n+1})$.
  • For the Hénon-Heiles potential
    $U(p,q)=p^{2}q-\frac{1}{3}q^{3},$
    the two-point Kahan discretized gradient (with $k=1$) is
    $\nabla_{\mathrm{K}}U(r_{n},r_{n+1})=\begin{pmatrix}p_{n}q_{n+1}+p_{n+1}q_{n}\\ p_{n}p_{n+1}-q_{n}q_{n+1}\end{pmatrix}\qquad r=(p,q)^{\mathsf{T}},$
    and the corresponding coefficient matrix and vector in (20) are $B_{n}=\begin{pmatrix}q_{n}&p_{n}\\ p_{n}&-q_{n}\end{pmatrix}$ and $\ell_{n}=0$.
Before presenting the properties of scheme (19), we give the following notations and assumptions:
Stacked state and history norm
Define the stacked state
$Y_{n}=(x_{n+k-1},x_{n+k-2},\dots,x_{n})\in(\mathbb{R}^{d})^{k},$
and the norm on the $k$-step history
$\|Y_{n}\|_{\infty}=\max_{0\leq j\leq k-1}\|x_{n+j}\|.$
$k$-step assumptions
For the $k$-step scheme (19), we assume that along the numerical trajectory the matrices $B_{n}$ in Lemma 3 satisfy $\|B_{n}\|\leq M_{B}$, and the uniform bound $M_{\phi}=\max\limits_{0\leq s\leq kT}\|\phi(sA)\|$ holds for a fix final time $T>0$. Finally, we assume the mappings $(x_{n},\ldots,x_{n+k-1})\mapsto B_{n}$ and $(x_{n},\ldots,x_{n+k-1})\mapsto\ell_{n}$ are Lipschitz on the relevant bounded history set.
Theorem 3.
Assuming the $k$-step assumptions, then the $k$-step EKahan scheme (19) is well-defined for a sufficiently small $h$ and stable on $[0,T]$ with respect to starting perturbations.
Proof.
By Lemma 3, (19) is equivalent to
$\bigl(I-kh\,\phi(khQM)\,QB_{n}\bigr)x_{n+k}=e^{khQM}x_{n}+kh\,\phi(khQM)\,Q\ell_{n}.$ (21)
Under the $k$-step assumptions, there will exist a $h_{0}$ such that
$kh\,M_{\phi}\,\|Q\|\,M_{B}\leq\theta<1\quad\text{for all }0<h\leq h_{0}.$ (22)
Then the coefficient matrix $I-kh\,\phi(khA)\,QB_{n}$ is invertible and Neumann-series bound further gives
$\bigl\|\bigl(I-kh\,\phi(khA)\,QB_{n}\bigr)^{-1}\bigr\|\leq\frac{1}{1-\theta}\quad\text{for all }n\text{ and }0<h\leq h_{0}.$ (23)
Therefore $x_{n+k}$ is uniquely determined.
Let $\{x_{n}\}$ and $\{\tilde{x}_{n}\}$ be two numerical solutions produced by (19) with the same step size. Set the error of the current state by $\delta_{n}=x_{n}-\tilde{x}_{n}$ and the error of the stacked state by $\Delta_{n}=Y_{n}-\tilde{Y}_{n}$ . Define
$G_{n}=kh\,\phi(khQM)\,QB_{n},\qquad b_{n}=e^{khQM}x_{n}+kh\,\phi(khQM)\,Q\ell_{n},$
and $\tilde{G}_{n},\tilde{b}_{n}$ analogously. Then from (21) we have $(I-G_{n})x_{n+k}=b_{n}$ and $(I-\tilde{G}_{n})\tilde{x}_{n+k}=\tilde{b}_{n}$. Subtracting and using the resolvent identity yields
$\delta_{n+k}=(I-G_{n})^{-1}(b_{n}-\tilde{b}_{n})+(I-G_{n})^{-1}(G_{n}-\tilde{G}_{n})(I-\tilde{G}_{n})^{-1}\tilde{b}_{n}.$ (24)
Taking $h_{0}$ such that $\theta=\frac{1}{2}$ in (22)-(23), we have $kh\,M_{\phi}\,\|Q\|\,M_{B}\leq\tfrac{1}{2}$ and
$\|(I-G_{n})^{-1}\|\leq\frac{1}{1-\|G_{n}\|}\leq 1+2\|G_{n}\|\leq 1+C_{B}h,\quad C_{B}:=2kM_{\phi}\|Q\|M_{B},$ (25)
and the same bound holds for $\|(I-\tilde{G}_{n})^{-1}\|$.
Moreover, recalling from the proof of Theorem 1, we have $\|e^{khQM}\|\leq 1+C_{E}h$ with $C_{E}=k\|QM\|M_{E}$. Taking norms in (24) and using (25) yield
$\|\delta_{n+k}\|\leq(1+C_{B}h)\,\|b_{n}-\tilde{b}_{n}\|+(1+C_{B}h)^{2}\,\|G_{n}-\tilde{G}_{n}\|\,\|\tilde{b}_{n}\|.$ (26)
Furthermore, we have
$\begin{split}\|b_{n}-\tilde{b}_{n}\|&\leq\|e^{khQM}\|\,\|\delta_{n}\|+kh\,\|\phi(khQM)\|\,\|Q\|\,\|\ell_{n}-\tilde{\ell}_{n}\|\\ &\leq(1+C_{E}h)\|\delta_{n}\|+kh\,M_{\phi}\|Q\|\,\|\ell_{n}-\tilde{\ell}_{n}\|,\end{split}$
and
$\|G_{n}-\tilde{G}_{n}\|\leq kh\,\|\phi(khQM)\|\,\|Q\|\,\|B_{n}-\tilde{B}_{n}\|\leq kh\,M_{\phi}\|Q\|\,\|B_{n}-\tilde{B}_{n}\|.$
By the Lipschitz assumption on the mappings $(x_{n},\ldots,x_{n+k-1})\mapsto B_{n}$ and $(x_{n},\ldots,x_{n+k-1})\mapsto\ell_{n}$, there exist constants $L_{B},L_{\ell}>0$ such that
$\|B_{n}-\tilde{B}_{n}\|\leq L_{B}\|\Delta_{n}\|_{\infty},\qquad\|\ell_{n}-\tilde{\ell}_{n}\|\leq L_{\ell}\|\Delta_{n}\|_{\infty}.$
In addition, boundedness of the trajectories implies $\|\tilde{b}_{n}\|\leq M_{b}$ for some constant $M_{b}>0$. Substituting these estimates into (26) and using $\|\delta_{n}\|\leq\|\Delta_{n}\|_{\infty}$ gives
$\|\delta_{n+k}\|\leq\Bigl((1+C_{B}h)(1+C_{E}h+khM_{\phi}\|Q\|L_{\ell})+kh(1+C_{B}h)^{2}M_{\phi}\|Q\|L_{B}\,M_{b}\Bigr)\|\Delta_{n}\|_{\infty}.$
Therefore, for $0<h\leq h_{0}$ there exists $\widetilde{C}>0$ independent of $h$ such that
$\|\delta_{n+k}\|\leq(1+\widetilde{C}h)\|\Delta_{n}\|_{\infty}.$
Consequently,
$\|\Delta_{n+1}\|_{\infty}=\max\bigl\{\|\delta_{n+k}\|,\ \|\delta_{n+k-1}\|,\ldots,\|\delta_{n+1}\|\bigr\}\leq(1+\widetilde{C}h)\|\Delta_{n}\|_{\infty}.$ (27)
Iterating and using $nh\leq T$ yields
$\|\Delta_{n}\|_{\infty}\leq(1+\widetilde{C}h)^{n}\|\Delta_{0}\|_{\infty}\leq e^{\widetilde{C}T}\|\Delta_{0}\|_{\infty},\qquad 0\leq nh\leq T,$
which completes the proof. ∎
Moreover, the properties of EKahan scheme established for systems with a cubic Hamiltonian can be extended to higher-degree polynomial Hamiltonians, as summarized next.
Theorem 4.
For semilinear system (2) with polynomial potential $U(x)$ of degree $k+2$, the EKahan scheme (19) is symmetric.
Proof.
From equation (19) we get
$x_{n+k}-e^{khA}x_{n}=kh\,\phi(khA)\,Q\,\nabla_{\mathrm{K}}U(x_{n},\ldots,x_{n+k}),\qquad A=QM.$ (28)
Using $\phi(-Z)=e^{-Z}\phi(Z)$ and $e^{-khA}=(e^{khA})^{-1}$, we can rewrite (28) as
$x_{n}-e^{-khA}x_{n+k}=-kh\,\phi(-khA)\,Q\,\nabla_{\mathrm{K}}U(x_{n},\ldots,x_{n+k}).$
Since $\nabla_{\mathrm{K}}U$ is symmetric in its arguments, $\nabla_{\mathrm{K}}U$ will not change if we reverse the sequence $(x_{n},\ldots,x_{n+k})$. Therefore, the method is symmetric. ∎
Lemma 4.
Let $A\in\mathbb{R}^{d\times d}$ and $g\in C^{2}([0,kh];\mathbb{R}^{d})$. For all $t\in[0,kh]$, we claim
$\int_{0}^{kh}e^{(kh-s)A}g(s)\mathrm{d}s=kh\,\phi(khA)\,g(kh/2)+R,\qquad\|R\|\leq Ch^{3},$
where $C$ depends on $k$, $M_{E}$, $\|A\|$, and $\max\limits_{0\leq s\leq kh}\bigl(\|g(s)\|+\|g^{\prime}(s)\|+\|g^{\prime\prime}(s)\|\bigr)$.
Lemma 5.
Assume $U$ is homogeneous and let $x(t)$ be a smooth solution on $t_{n},t_{n}+kh]$. Set $\tilde{x}_{n+i}:=x(t_{n}+ih)$ and $t_{n+\frac{k}{2}}:=t_{n}+\frac{k}{2}h$. Then there exists $C>0$, independent of $h$, such that
$\bigl\|\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k})-\nabla U\bigl(x(t_{n+\frac{k}{2}})\bigr)\bigr\|\leq Ch^{2}.$
Lemma 6.
Let $x(t)$ solve (2) with $U$ homogeneous of degree $k+2$ and set $\tilde{x}_{n+i}=x(t_{n}+ih)$. Then the local defect of (19) satisfies
$\|\tau_{n+k}\|\leq C_{\mathrm{loc}}\,h^{3},\qquad t_{n}\leq T,$
for a constant $C_{\mathrm{loc}}$ independent of $h$.
Proof.
By the variation of constants formula for $\dot{x}=Ax+Q\nabla U(x)$, we have
$\tilde{x}_{n+k}=e^{khA}\tilde{x}_{n}+\int_{0}^{kh}e^{(kh-s)A}Q\nabla U(x(t_{n}+s))\mathrm{d}s.$
Hence
$\tau_{n+k}=\int_{0}^{kh}e^{(kh-s)A}Q\nabla U(x(t_{n}+s))\mathrm{d}s-kh\,\phi(khA)\,Q\,\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k}).$
Add and subtract the exponential midpoint quantity $kh\,\phi(khA)\,Q\,\nabla U(x(t_{n+\frac{k}{2}}))$:
$\begin{split}\tau_{n+k}&=\underbrace{\Bigl[\int_{0}^{kh}e^{(kh-s)A}Q\nabla U(x(t_{n}+s))\mathrm{d}s-kh\,\phi(khA)\,Q\,\nabla U(x(t_{n+\frac{k}{2}}))\Bigr]}_{=:I_{1}}\\ &+\underbrace{kh\,\phi(khA)\,Q\Bigl[\nabla U(x(t_{n+\frac{k}{2}}))-\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k})\Bigr]}_{=:I_{2}}.\end{split}$
For $I_{1}$, apply Lemma 4 with $g(s)=Q\nabla U(x(t_{n}+s))$ to get $\|I_{1}\|\leq Ch^{3}$. For $I_{2}$, use the boundedness of $\phi(khA)$ on $[0,T]$ and Lemma 5 to obtain
$\|I_{2}\|\leq kh\,\|\phi(khA)\|\,\|Q\|\,\bigl\|\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k})-\nabla U(x(t_{n+\frac{k}{2}}))\bigr\|\leq C\,h\cdot h^{2}=Ch^{3}.$
Combining the two bounds yields $\|\tau_{n+k}\|\leq C_{\mathrm{loc}}h^{3}$. ∎
Lemma 7.
Assume the exact solution stays in a bounded set $\mathcal{D}$ on $[0,T]$ and $U$ is polynomial. Then there exist $h_{0}>0$ and $C>0$ such that for all $0<h\leq h_{0}$ the $k$-step EKahan update defines a unique one-step map $Y_{n+1}=\Phi_{h}(Y_{n})$ on $\mathcal{D}^{k}$, and
$\|\Phi_{h}(Y)-\Phi_{h}(Z)\|_{\infty}\leq(1+Ch)\,\|Y-Z\|_{\infty},\qquad\forall\,Y,Z\in\mathcal{D}^{k},$
Theorem 5.
Under the assumptions of Lemma 7, suppose the starting values $x_{0},\ldots,x_{k-1}$ satisfy $\|x_{j}-\tilde{x}_{j}\|\leq Ch^{2}$ for $j=0,\ldots,k-1$. Then for $t_{n}=nh\leq T$ and $h\leq h_{0}$,
$\max_{0\leq n\leq N}\|x_{n}-\tilde{x}_{n}\|\leq C_{T}\,h^{2},$
i.e., the $k$-step EKahan scheme is globally second order on finite time intervals.
Proof.
Inserting the exact solution values $\tilde{Y}_{n}$ into the scheme gives
$\tilde{Y}_{n+1}=\Phi_{h}(\tilde{Y}_{n})+E_{n+1},\qquad\|E_{n+1}\|\leq C_{\mathrm{loc}}h^{3}.$
where $E_{n}$ contains only the last component defect $\tau_{n+k}$ from Lemma 6. Subtracting from $Y_{n+1}=\Phi_{h}(Y_{n})$ and using Lemma 7 yields the following recursion for the extended error $\Delta_{n}=Y_{n}-\tilde{Y}_{n}$
$\|\Delta_{n+1}\|_{\infty}\leq(1+Ch)\|\Delta_{n}\|_{\infty}+C_{\mathrm{loc}}h^{3}.$
Applying the discrete Gronwall inequality gives, for $t_{n}\leq T$,
$\|\Delta_{n}\|_{\infty}\leq e^{CT}\|\Delta_{0}\|_{\infty}+\frac{C_{\mathrm{loc}}}{C}\,(e^{CT}-1)\,h^{2}.$
Since $\|\Delta_{0}\|_{\infty}=\max\limits_{0\leq j\leq k-1}\|x_{j}-\tilde{x}_{j}\|=O(h^{2})$ by assumption, the claim follows. ∎